OpenStudy (anonymous):

the cost of 3 doughnuts and 5 cups of coffee is 9.90 but 2 doughnuts and 3 coffees cost 6.15 Find the cost of 1 doughnut and 1 cup of coffee. )equations/solution required)

OpenStudy (lgbasallote):

let x be doughnuts and y be coffee so.. 3x + 5y = 9.90 and 2x + 3y = 6.15 make sense?

OpenStudy (anonymous):

Dang, I could go for some doughnuts and coffee right now . . .

OpenStudy (anonymous):

i don't know how to solve the equation. help me

OpenStudy (anonymous):

It's a system of equations: 2 equations with 2 unknowns. You'll have to substitute one into the other to do it.

OpenStudy (anonymous):

i don't know how to it

OpenStudy (anonymous):

OK, I'm going to multiply everything by 100 cents/$ to get rid of the decimals (I don't like decimals). I'm going to call doughnuts, 'D' and coffee, 'C'. 300D+500C=990 200D+300C=615

OpenStudy (anonymous):

Solve the first equation for C in terms of D and solve the second equation for D in terms of C.

OpenStudy (anonymous):

Can you do that much? After that, substitute the expression for C into the equation for D.

OpenStudy (anonymous):

sorry i just have a learning disability so I'm struggling.

OpenStudy (anonymous):

Should look like this: \[C=\frac{990-300D}{500} \space , \space D=\frac{615-300C}{200}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You can substitute C into D to get a number for D, then put that number back into the equation for C to get a number for C.